OpenStudy (anonymous):

The value of a professional basketball player's autograph rose 20% in the last year. It is now worth $288.00. What was it worth a year ago? $220.00 $240.00 $230.00 $260.00

OpenStudy (anonymous):

@hlee13

OpenStudy (anonymous):

Sorry .. I don't do math .. I'm strickly history and government .. everything else I wouldn't be comfortable

OpenStudy (texaschic101):

x + .20x = 288 1.20x = 288 x = 288/1.20 x = 240 check... 240 + .20(240) = 288 240 + 48 = 288 288 = 288 (correct) answer is 240

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

A zookeeper predicted the weight of a new baby elephant to be 258 pounds when it was born. The elephant actually weighed 290 pounds at birth. What was the percent error of the zookeeper's prediction? 88.97% 12.40% 11.03% 32%

OpenStudy (texaschic101):

290 - 258 = 32 32/290 = 0.1103 changed to percent = 11.03% I think that is how you do it. You might better get a second opinion

OpenStudy (anonymous):

Sam opened a money-market account that pays 2% simple interest. He started the account with $7,000 and made no further deposits. When he closed the account, he had earned $420 in interest. How long did he keep his account open? 4 years 3 years 6 years 5 years

OpenStudy (anonymous):

that was right thz

OpenStudy (texaschic101):

I = PRT 420 = (7000)(.02)(T) 420 = 140T 420/140 = T 3 = T It took 3 years

OpenStudy (anonymous):

3. Bert is planning to open a savings account that earns 1.6% simple interest yearly. He wants to earn exactly $256 in interest after 2 years. How much money should he deposit? $800 $80 $8,000 $4,096

OpenStudy (texaschic101):

I = PRT 256 = P(.016)(2) 256 = .032P 256/.032 = P 8000 = P I think I did that right...answer is 8000

OpenStudy (anonymous):

Andy estimated that he would need 72 feet of lumber for a tree house project. He later found that the actual amount of lumber needed was 61 feet. What was the percent error of Andy's estimation? 15.28% 11% 84.72% 18.03%

OpenStudy (anonymous):

@texaschic101